OpenStudy (anonymous):

The length of a rectangle is 5mm longer than its width. Its perimeter is 30mm. let w equal the width of the rectangle.

OpenStudy (anonymous):

2(W + L) = 30 2(W + W + 5) = 30 W = 5 L = 10

OpenStudy (johnweldon1993):

Okay well we know the total perimeter of a rectangle is 2L + 2W = P we know P = 30 so 2L + 2W = 30 now we are also given that the length of the rectangle is 5mm longger than the width So L = W + 5 so 2L + 2W = 30 L = W + 5 2 equations...2 unknowns....can you solve this?

OpenStudy (anonymous):

still need help !!

OpenStudy (johnweldon1993):

Alright...well if you look at the equation equation...we see that L = W + 5 Well since we know that 'L' equals THAT...we can substitute that into the first equation... So everytime you see L in the first equation...replace it with (W + 5) 2L + 2W = 30 would become 2(W + 5) + 2W = 30 Distribute out the 2... 2W + 10 + 2W = 30 Can you go on from here?